find the solution to problem by using system of equations by substitution.
y=2x+7
y=-x-5

Answer by rfer(2688)

y=2x+7
y=-x-5
---------------
-x-5=2x+7
3x=-12
x=-4
----------------
y=2(-4)+7
y=-8+7
y=-1
-------------
-1=4-5